# ProSparse: Introducing and Enhancing Intrinsic Activation Sparsity within Large Language Models

### v2

@@ Abstract
Activation sparsity refers to the existence of considerable weakly-contributed elements among activation outputs. As a prevalent property of the models using the ReLU activation function, it has been proven a promising paradigm to boost model inference efficiency. Nevertheless, most large language models (LLMs) adopt activation functions without intrinsic activation sparsity (e.g., GELU and Swish). Some recent efforts have explored introducing ReLU or its variants as the substitutive activation function to help LLMs achieve activation sparsity and inference acceleration, but few can simultaneously obtain high sparsity and comparable model performance. This paper introduces an effective sparsification method named “ProSparse” to push LLMs for higher activation sparsity without decreasing model performance. Specifically, after substituting the activation function of LLMs with ReLU, ProSparse adopts progressive sparsity regularization with a factor smoothly increasing along sine curves in multiple stages. This can enhance activation sparsity and alleviate performance degradation by avoiding radical shifts in activation distribution. With ProSparse, we obtain high sparsity of 89.32% and 88.80% for LLaMA2-7B and LLaMA2-13B, respectively, achieving comparable performance to their original Swish-activated versions. Our inference acceleration experiments further demonstrate the practical acceleration brought by higher activation sparsity.
@@ 1 Introduction
Recent years have witnessed the significant breakthrough made by large language models (LLMs), and these LLMs display commendable performance across a wide range of NLP tasks Brown et al. (2020); Wei et al. (2021); Ouyang et al. (2022); OpenAI (2023); Touvron et al. (2023a, b); Achiam et al. (2023). Nevertheless, the formidable computational costs required by LLM deployment and inference pose a considerable challenge to the wider application of LLMs Aminabadi et al. (2022); Pope et al. (2023). Among various techniques for handling this challenge, the utilization of activation sparsity is a promising one, for its effectiveness in enhancing inference efficiency Liu et al. (2023); Song et al. (2023) by leveraging the weakly-contributed elements among the outputs of LLM activation functions.
Using ReLU, which naturally outputs zero elements, as the activation function is a straightforward method to achieve intrinsic activation sparsity and widely adopted in early LLMs Raffel et al. (2020); Zhang et al. (2022a). However, recent LLMs predominantly favor GELU and Swish Touvron et al. (2023a); Chowdhery et al. (2023); Almazrouei et al. (2023), and thus lack intrinsic activation sparsity. To pursue the sparsity-based inference acceleration, ReLUfication Zhang et al. (2022b); Mirzadeh et al. (2023) has been explored to introduce the ReLU-based intrinsic activation sparsity into non-ReLU LLMs. Preliminary ReLUfication methods Zhang et al. (2022b, 2024) directly substitute the activation functions of non-ReLU LLMs with ReLU. Since activation function substitution cannot overcome the inherent limitation imposed by the original non-ReLU activation distribution, Mirzadeh et al. (2023) further introduce the inserted and shifted ReLU functions to enforce higher sparsity through radically shifting the activation distribution. Despite the promise of ReLUfication, existing efforts fail to achieve satisfactory activation sparsity and risk performance degradation caused by ReLUfication.
@@
In this paper, we propose an effective progressive ReLUfication method to help non-ReLU LLMs obtain high activation sparsity without compromising performance. We name the method “ProSparse”, which includes three steps shown in Figure 1: activation function substitution, progressive sparsity regularization, and activation threshold shifting. The first step is to replace the activation function of non-ReLU LLMs with ReLU and then apply continual training for adapting LLM to the new ReLU activation. As discussed above, this can hardly achieve satisfactory sparsity. Therefore, in the second step, we apply sparsity regularization Ma et al. (2019) to the intermediate activation outputs of the feed-forward networks (FFNs) within LLMs to seek higher activation sparsity. Considering the potential performance risks of forcing the fixed regularization factor Ma et al. (2019); Li et al. (2020), we progressively increase the regularization factor in multiple stages. Concretely, the factor is set to a low constant value for the warmup stage. Next, during each subsequent stage, the factor undergoes a gradual increase along a gentle sine curve. Following the cosine annealing learning rate scheduler Loshchilov and Hutter (2016), such progressive sparsity regularization can provide more time for the model to adapt to increasing regularization and avoid a radical shift in activation distribution, thereby alleviating performance degradation. The final step, inspired by FATReLU Kurtz et al. (2020), modifies the vanilla ReLU by shifting its activation threshold to a positive value. This prunes less influential neurons to further improve sparsity.
@@
In our experiments, we apply ProSparse to the ReLUfication of LLaMA2 Touvron et al. (2023b). Activation sparsity of 89.32% and 88.80% are successfully achieved for LLaMA2-7B and LLaMA2-13B respectively, with performance comparable to original Swish-activated versions on various LLM benchmarks. Furthermore, we demonstrate the practical inference acceleration effect of higher activation sparsity by applying an approximate algorithm and an accurate algorithm to the inference of models with different sparsity. For the approximate one, we use PowerInfer Song et al. (2023), which achieves state-of-the-art speedup ratios tailored for sparsely-activated LLMs at the expense of potentially inaccurate inference due to the mistakes of activation predictors. For the accurate one, we implement two GPU operators that leverage the input-side and output-side sparsity during the computation of ReLU-activated FFN layers. [^1] Codes for these two operators are available at https://github.com/Raincleared-Song/sparse_gpu_operator. The experimental results demonstrate that those models with higher sparsity can achieve more significant inference acceleration with both algorithms.
@@
In summary, we make the following contributions in this paper: (1) We propose ProSparse, an effective ReLUfication method that converts non-ReLU LLMs into much sparser ReLU-activated LLMs without decreasing model performance. (2) Sparsely-activated versions of LLaMA2-7B and LLaMA2-13B comparable to their original Swish-activated versions in performance are both obtained and available. [^2] Available at https://huggingface.co/SparseLLM/prosparse-llama-2-7b and https://huggingface.co/SparseLLM/prosparse-llama-2-13b respectively. (3) We demonstrate the practical inference acceleration effect of higher activation sparsity that ProSparse can reach.
2 Related Works
@@
Here we mainly introduce works on improving LLM inference efficiency. More details on LLMs can refer to the existing surveys Bommasani et al. (2021); Zhao et al. (2023). More related works about $L_{1}$ regularization are listed in Appendix A.
@@
Despite the commendable performance of LLMs, the sustainable increase in the scale of model parameters brings the exponential growth of inference computations, making the deployment of LLMs a formidable challenge Kaplan et al. (2020); Liu et al. (2023). To reduce the computational costs required by the inference of such large models, various model compression methods have been proposed, such as quantization Han et al. (2015a); Jacob et al. (2018); Nagel et al. (2019); Zhao et al. (2019); Bai et al. (2022); Xiao et al. (2023); Yao et al. (2023), pruning Han et al. (2015a, b); Molchanov et al. (2016); Hoefler et al. (2021); Ma et al. (2023); Sun et al. (2023); Frantar and Alistarh (2023); Xia et al. (2023), and distillation Hinton et al. (2015); Tang et al. (2019); Touvron et al. (2021); Gu et al. (2023); Hsieh et al. (2023). Efficient sampling methods have also been proposed to achieve faster inference decoding speed Leviathan et al. (2023); Wang et al. (2023); Chen et al. (2023); Miao et al. (2023).
@@
In general, none of the above methods involve leveraging the intrinsic mechanisms within LLMs to achieve inference acceleration. To this end, some recent works Li et al. (2022); Liu et al. (2023); Song et al. (2023) notice the intrinsic activation sparsity within some LLMs and exploit this sparsity for inference acceleration. Activation sparsity refers to the phenomenon where certain model parameters, corresponding to those zero or small elements in activation outputs, have a weak impact on final LLM outputs given a specific input. These weakly-contributed parameters are regarded as inactivated and can thus be skipped during inference to achieve sparse computation, without redundant computational resources spent on them. Therefore, the utilization of activation sparsity is orthogonal to model compression and efficient sampling, and these approaches can be easily accumulated.
@@
Previous studies have marked activation sparsity as a prevalent phenomenon existing in almost any ReLU-activated Transformer architecture Li et al. (2022), from LLMs (e.g., T5 Raffel et al. (2020) and OPT Zhang et al. (2022a)) to some vision models (e.g., ViT Dosovitskiy et al. (2020)). However, recent LLMs such as Falcon Almazrouei et al. (2023) and LLaMa Touvron et al. (2023b) prevalently adopt non-ReLU activation functions such as GELU Hendrycks and Gimpel (2016) and Swish Elfwing et al. (2018) and do not exhibit activation sparsity. Therefore, to leverage the merits of activation sparsity without training a ReLU-activated LLM from scratch, many works conduct ReLUfication, which introduces sparse ReLU-based activations into non-ReLU LLMs. Zhang et al. (2022b) train the GELU-activated BERT Devlin et al. (2018) into a ReLU-activated version after a direct substitution of the activation function. Zhang et al. (2024) apply a similar paradigm to Falcon and LLaMA, which are activated by GELU and Swish respectively. Since activation substitution cannot reach satisfactory sparsity, mainly due to the unhandled inherent limitation of the original non-ReLU activation distribution, Mirzadeh et al. (2023) introduce the inserted and shifted ReLU activation functions and conduct a radical shift in activation distribution. Although these shifted operations are claimed to achieve sparsity of nearly 95%, we cannot replicate the results in our experiments (Section 4.2) and the sparsity is still limited.
@@
As discussed above, we can clearly recognize the promise of activation sparsity and also observe the key challenge of leveraging ReLUfication to achieve activation sparsity: how to concurrently achieve high sparsity and comparable performance. To this end, this paper introduces ProSparse, a ReLUfication method that can obtain high ReLU-based activation sparsity for non-ReLU LLMs without performance degradation.
3 Methodology
@@ 3.1 Definitions and Notations
For the convenience of subsequent demonstrations, here we define activation sparsity within LLMs in detail. Since the activation function mainly exists in the FFN layers of LLMs, we first formalize the computation process of FFNs. Given the hidden dimension $d_{model}$ and the FFN intermediate dimension $d_{ff}$, the computation process of a gated FFN (i.e., the most widely adopted FFN architecture in recent Transformer-based LLMs Dauphin et al. (2017); Shazeer (2020)) can be formalized as:
$\mathbf{s}=\sigma(\mathbf{x}\mathbf{W}_{s}^{T}),$ $\quad\mathbf{x}_{1}=\mathbf{s}\odot(\mathbf{x}\mathbf{W}_{1}^{T}),$ (1)
$\text{FFN}(\mathbf{x})$ $=\mathbf{x}_{1}\mathbf{W}_{2}^{T},$
@@
We define the activation sparsity (hereinafter abbreviated as sparsity) as the ratio of zero elements (i.e., inactivated elements) in $\mathbf{x}_{1}$ for a specific input $\mathbf{x}$. However, since the sparsity often varies in different layers for different inputs, we evaluate the sparsity of the whole LLM using the average sparsity, defined as the average value of sparsity across all layers in an LLM on a sufficiently large amount of input data.
@@ 3.2 Methodology of ProSparse
Nevertheless, activation function substitution by nature does not change the activation distribution, which will potentially limit the sparsity to relatively low values. To push for higher sparsity, a typical method is $L_{1}$ sparsity regularization Li et al. (2022), which introduces the $L_{1}$ regularization loss as an extra training target. Given the intermediate output $\mathbf{x}_{1}$ of the $i$-th FFN layer in an LLM, the regularization loss is defined as:
$\mathcal{L}_{reg}^{i}=\lambda\cdot||\mathbf{x}_{1}||_{1},$ (2)
where $||\cdot||_{1}$ is the $L_{1}$ norm operator and $\lambda$ is the regularization factor. For an LLM with $K$ FFN layers, the total regularization loss is summed from the losses of all layers, namely $\mathcal{L}_{reg}=\sum_{i=1}^{K}\mathcal{L}_{reg}^{i}$.
Algorithm 1 The process of progressive sparsity regularization in ProSparse.
@@
0:  Accumulated step numbers of respective stages $\{T_{i}\}_{i=1}^{S}$, s.t. $0<T_{1}<T_{2}<...<T_{S}$.
@@
4:  end for
5:  // incremental stages
6:  for $i\leftarrow 2$ to $S$ do
7:     for $t\leftarrow T_{i-1}+1$ to $T_{i}$ do
8:        $\eta\leftarrow\frac{1}{2}[\sin(-\frac{\pi}{2}+\frac{t-T_{i-1}}{T_{i}-T_{i-1}}\pi)+1]$
9:        $\lambda\leftarrow\lambda_{i-1}+\eta(\lambda_{i}-\lambda_{i-1})$
10:     end for
11:  end for
Considering the potentially unstable training and performance degradation due to fixed regularization factors Georgiadis (2019); Kurtz et al. (2020); Li et al. (2022), we propose the progressive sparsity regularization, where the factor $\lambda$ is carefully scheduled to gently increase in multiple stages, as displayed in Algorithm 1.
Concretely, for the warmup stage, we set $\lambda$ to a constant value, which is relatively small to prevent radical activation distribution shifts and introduce higher preliminary sparsity. Next, for each of the remaining stages (hereinafter called incremental stages), $\lambda$ is scheduled to increase along a smooth sine curve from a trough value to a peak value. Inspired by the cosine annealing scheduler for learning rates Loshchilov and Hutter (2016), we choose the sine function for $\lambda$ increase owing to its special trend. Specifically, the sine function has small derivatives near the trough and the peak, thereby $\lambda$ will not increase radically around these two points. This provides the LLMs with more time to adapt the activation distributions to the newly increased $L_{1}$ regularization. Notably, each stage is accompanied by certain steps of training. The step number and peak value of each stage are chosen according to the target sparsity and stability requirements.
@@
As demonstrated by recent works, there exist considerable amounts of non-zero low elements in the activation outputs, which have little influence on final results and thus can be pruned for higher sparsity Zhang et al. (2024). Therefore, following FATReLU Kurtz et al. (2020), we modify the ReLU by shifting the activation threshold, i.e.,
@@
where $T>0$ is a positive threshold. As long as $T$ is properly chosen (see Appendix G), such an adjustment can increase sparsity with negligible losses Zhang et al. (2024).
3.3 Practical Inference Acceleration Test
To go further beyond former theoretical acceleration analyses based on FLOPs (Floating Point Operations Per Second) Mirzadeh et al. (2023) and establish the practical value of ProSparse, we compare the acceleration effects of LLMs with different sparsity on real GPU hardware. For comprehensiveness, we introduce two categories of acceleration algorithms based on activation sparsity: an approximate algorithm and an accurate algorithm.

### v3

@@ Abstract
Activation sparsity refers to the existence of considerable weakly-contributed elements among activation outputs. As a prevalent property of the models using the ReLU activation function, activation sparsity has been proven a promising paradigm to boost model inference efficiency. Nevertheless, most large language models (LLMs) adopt activation functions without intrinsic activation sparsity (e.g., GELU and Swish). Some recent efforts have explored introducing ReLU or its variants as the substitutive activation function to help LLMs achieve activation sparsity and inference acceleration, but few can simultaneously obtain high sparsity and comparable model performance. This paper introduces a simple and effective sparsification method named “ProSparse” to push LLMs for higher activation sparsity while maintaining comparable performance. Specifically, after substituting the activation function of LLMs with ReLU, ProSparse adopts progressive sparsity regularization with a factor smoothly increasing along the multi-stage sine curves. This can enhance activation sparsity and mitigate performance degradation by avoiding radical shifts in activation distributions. With ProSparse, we obtain high sparsity of 89.32% for LLaMA2-7B, 88.80% for LLaMA2-13B, and 87.89% for end-size MiniCPM-1B, respectively, achieving comparable performance to their original Swish-activated versions. These present the most sparsely activated models among open-source LLaMA versions and competitive end-size models, considerably surpassing ReluLLaMA-7B (66.98%) and ReluLLaMA-13B (71.56%). Our inference acceleration experiments further demonstrate the significant practical acceleration potential of LLMs with higher activation sparsity, obtaining up to 4.52$\times$ inference speedup.
@@ 1 Introduction
Recent years have witnessed the significant breakthrough made by large language models (LLMs), and these LLMs display commendable performance across a wide range of NLP tasks [8, 77, 54, 53, 72, 73, 1]. Nevertheless, the formidable computational costs required by the deployment and inference of LLMs pose a considerable challenge to the wider application of LLMs [3, 56]. The utilization of activation sparsity is one of the most promising techniques to reduce computational costs and enhance inference efficiency [44, 66], by discarding the redundant computation associated with the elements among LLM activation outputs that contribute weakly to the final result.
The adoption of ReLU, which naturally outputs zero elements, as the activation function is a straightforward method to achieve intrinsic activation sparsity and widely adopted in early LLMs [58, 86]. However, recent LLMs predominantly favor non-ReLU activation functions, such as GELU and Swish [72, 12, 2]. Although these non-ReLU LLMs may also display activation sparsity according to recent findings [88], such sparsity is manually imposed by searching adaptive activation thresholds (i.e., non-intrinsic), which can potentially lose minor neuron outputs and degrade performance. To pursue the sparsity-based inference acceleration, the task of ReLUfication is proposed [87, 50], aiming to introduce ReLU-based intrinsic activation sparsity into non-ReLU LLMs. Preliminary ReLUfication methods [87, 88] directly substitute the activation functions of non-ReLU LLMs with ReLU. As such substitution cannot overcome the inherent limitation imposed by the original non-ReLU activation distribution, inserted and shifted ReLU functions [50] are further introduced to enforce higher sparsity through radically shifting the activation distribution. Despite the promise of ReLUfication, existing efforts fail to achieve satisfactory activation sparsity and risk performance degradation caused by ReLUfication.
@@
In this paper, we propose a simple and effective ReLUfication method to help non-ReLU LLMs obtain high activation sparsity without performance degradation. We name the method “ProSparse”, which includes three steps shown in Figure 1: activation function substitution, progressive sparsity regularization, and activation threshold shifting. The first step is to replace the activation function of non-ReLU LLMs with ReLU and then apply continual training for adapting LLMs to the new ReLU activation. As discussed above, this can hardly achieve satisfactory sparsity. Therefore, in the second step, we apply sparsity regularization [47] to the intermediate activation outputs of the feed-forward networks (FFNs) within LLMs to seek higher activation sparsity. Considering the potential performance risks of forcing the fixed regularization factor [47, 41], we progressively increase the regularization factor in multiple stages. Concretely, the factor is set to a low constant value for the warmup stage. Next, during each subsequent stage, the factor undergoes a gradual increase along a gentle sine curve. Such progressive sparsity regularization can provide more time for the model to adapt to increasing regularization and avoid a radical shift in activation distribution, thereby mitigating performance degradation. The final step adopts FATReLU [38] to modify the vanilla ReLU by shifting its activation threshold to a positive value. This prunes less influential neurons to further improve sparsity.
@@
In experiments, we apply ProSparse to the ReLUfication of LLaMA2 [73] and end-size MiniCPM [34]. Activation sparsity of 89.32%, 88.80%, and 87.89% are successfully achieved for LLaMA2-7B, LLaMA2-13B, and MiniCPM-1B, respectively, with performance comparable to their original Swish-activated versions on various LLM benchmarks. Furthermore, we demonstrate the practical inference acceleration effect of higher activation sparsity, by respectively applying an approximate algorithm and an accurate algorithm to the inference of models with different sparsity. For the approximate one, we use PowerInfer [66], which achieves state-of-the-art speedup ratios tailored for sparsely activated LLMs at the expense of potentially inaccurate inference due to the mistakes of activation predictors. For the accurate one, we implement and release two GPU operators that leverage the input-side and output-side sparsity during the computation of ReLU-activated FFN layers [^1] Source codes for these two operators are available at https://github.com/Raincleared-Song/sparse_gpu_operator..
The experimental results demonstrate that models with higher sparsity can achieve more significant inference acceleration with both approximate and accurate algorithms (e.g., up to 4.52$\times$ speedup with PowerInfer). Moreover, comprehensive analyses are conducted to figure out the quantitative relationship between the activation sparsity and the regularization factor, making the activation sparsity obtained by ProSparse more controllable. We also discuss the rationality of progressive $L_{1}$ regularization, empirical methods of performing supervised fine-tuning (SFT) on sparsely activated models, and the sparsity distribution among distinct datasets or layers.
@@
In summary, we make the following contributions in this paper: (1) We propose ProSparse, an effective ReLUfication method that converts non-ReLU LLMs into much sparser ReLU-activated LLMs without performance degradation. (2) Sparsely activated versions of LLaMA2-7B, LLaMA2-13B, and MiniCPM-1B comparable to their original Swish-activated versions in performance are both obtained and available [^2] Models are available at https://huggingface.co/SparseLLM/prosparse-llama-2-7b, https://huggingface.co/SparseLLM/prosparse-llama-2-13b, and https://huggingface.co/SparseLLM/ProSparse-MiniCPM-1B-sft, respectively.. (3) We demonstrate the practical inference acceleration effect of higher activation sparsity that ProSparse can reach. Valuable observations and analyses are also conducted.
2 Preliminaries and Related Works
@@
Here we mainly introduce works on improving LLM inference efficiency. Refer to existing surveys [7, 91] for more details about LLMs. Related works about $L_{1}$ regularization are listed in Appendix A.
Inference Acceleration for LLMs
@@
Despite the commendable performance of LLMs, the sustainable increase in model scales brings the exponential growth of inference computations, making the deployment of LLMs a formidable challenge [37, 44]. To reduce the computational costs required by LLM inference, various model compression methods have been proposed, such as quantization [25, 35, 52, 90, 5, 82, 83], pruning [25, 26, 51, 31, 48, 67, 21, 81], and distillation [30, 69, 71, 24, 33]. Efficient sampling methods have also been proposed to achieve faster inference decoding speed [39, 75, 9, 49]. In general, none of these acceleration methods leverages the intrinsic mechanisms within LLMs.
Activation Sparsity
@@
Recent works [43, 44, 66] have noticed the intrinsic activation sparsity within some LLMs and its potential in inference acceleration. Activation sparsity refers to the phenomenon where considerable zero or negligible elements in activation outputs, corresponding to certain model parameters (i.e., neurons), have a weak impact on LLM outputs given a specific input. These weakly-contributed parameters are regarded as inactivated and can thus be skipped during inference to save considerable computational resources. Notably, the utilization of activation sparsity is orthogonal to model compression and efficient sampling, and these approaches can be readily accumulated.
Difference between Activation Sparsity and Pruning
Generally, pruning realizes sparsity by removing certain elements (e.g., neurons, weights, or structured blocks) in LLMs. However, the sparsity introduced by pruning is statically limited to model weights and independent of the inputs. High static sparsity is often accompanied by considerable performance degradation compared to the original dense model [21, 81]. By contrast, activation sparsity is dynamically determined by the input data and thus potentially compromises less model capacity and downstream task performance.
ReLUfication
@@
Previous studies have marked the existence of activation sparsity in almost any ReLU-activated Transformer architecture [43], from LLMs (e.g., T5 [58] and OPT [86]) to vision models (e.g., ViT [19]). However, recent LLMs such as Falcon [2] and LLaMA [73] prevalently adopt non-ReLU activation functions such as GELU [29] and Swish [20] without intrinsic activation sparsity. Therefore, to leverage the merits of activation sparsity without training a ReLU-activated LLM from scratch, many works conduct ReLUfication, the task of introducing sparse ReLU-based activations into non-ReLU LLMs. In MoEfication [87], a GELU-activated BERT [17] is converted into a ReLU-activated version through direct activation function substitution and additional training. ReluLLaMA and ReluFalcon apply a similar paradigm to Falcon and LLaMA, respectively [88]. Since activation substitution cannot reach satisfactory sparsity, mainly due to the unhandled limitation of the original non-ReLU activation distribution, the inserted and shifted ReLU activation functions are introduced [50], conducting a radical shift in activation distribution. Although these operations are claimed to achieve sparsity of nearly 95%, we cannot replicate the results in our experiments (see the 3rd paragraph of Section 4.4) and the sparsity is still limited.
@@
As discussed above, we can clearly recognize the promise of activation sparsity and also observe the key challenge of leveraging ReLUfication to achieve activation sparsity: how to achieve high sparsity and mitigate performance degradation concurrently. These are what ProSparse is designed for.
3 Methods
@@ 3.1 Definitions and Notations
For the convenience of subsequent demonstrations, here we define activation sparsity within LLMs in detail. Since the activation function mainly exists in the FFN layers of LLMs, we first formalize the computation process of FFNs. Given the hidden dimension $d_{model}$ and the FFN intermediate dimension $d_{ff}$, the computation process of a gated FFN (i.e., the most widely adopted FFN architecture in recent Transformer-based LLMs [16, 64]) can be formalized as:
$\mathbf{s}=\sigma(\mathbf{x}\mathbf{W}_{s}^{T}),\quad\mathbf{x}_{1}=\mathbf{s}\odot(\mathbf{x}\mathbf{W}_{1}^{T}),\quad\text{FFN}(\mathbf{x})=\mathbf{x}_{1}\mathbf{W}_{2}^{T},$ (1)
@@
We define the activation sparsity (hereinafter abbreviated as sparsity) as the ratio of zero elements (i.e., inactivated elements) in $\mathbf{x}_{1}$ for a specific input $\mathbf{x}$. Since the sparsity varies in different layers for different inputs, we evaluate the sparsity of an LLM using the average sparsity, defined as the average value of sparsity across all layers in an LLM on a sufficiently large amount of input data.
@@ 3.2 ProSparse
Nevertheless, activation function substitution by nature does not change the activation distribution, which will potentially limit the sparsity to relatively low values. To push for higher sparsity, a typical method is $L_{1}$ sparsity regularization [43], which introduces the $L_{1}$ regularization loss as an extra training target. Given the intermediate output $\mathbf{x}_{1}$ of the $i$-th FFN layer in an LLM, the regularization loss is defined as:
$\mathcal{L}_{reg}^{i}(\lambda)=\lambda\cdot||\mathbf{x}_{1}||_{1},$ (2)
where $||\cdot||_{1}$ is the $L_{1}$ norm operator and $\lambda$ is the regularization factor. For an LLM with $K$ FFN layers, the total regularization loss is summed from the losses of all layers, namely $\mathcal{L}_{reg}(\lambda)=\sum_{i=1}^{K}\mathcal{L}_{reg}^{i}(\lambda)$. The overall optimization target is $\mathcal{L}_{lm}+\mathcal{L}_{reg}(\lambda)$, where $\mathcal{L}_{lm}$ is the vanilla language modeling loss.
Considering the potentially unstable training and performance degradation due to fixed regularization factors [23, 38, 43], we propose the progressive sparsity regularization, where the factor $\lambda$ is carefully scheduled to gently increase in multiple stages, as displayed in Algorithm 3.2.
Concretely, for the warmup stage, we set $\lambda$ to a constant value, which is relatively small to prevent radical activation distribution shifts and introduce higher preliminary sparsity. Next, for each of the remaining stages (called incremental stages), $\lambda$ is scheduled to increase along a smooth sine curve from a trough value to a peak value. Inspired by the cosine annealing scheduler for learning rates [46], we choose the sine function for $\lambda$ increase owing to its special trend. Specifically, the sine function has small derivatives near the trough and the peak, thereby $\lambda$ will not increase radically around these two points. This provides the LLMs with more time to adapt the activation distributions to the newly increased $L_{1}$ regularization. Notably, each stage is accompanied by certain steps of training. The step numbers and peak $\lambda$ values are chosen according to the target sparsity and stability requirements.
{wrapfloat}
algorithm[18]r0.5 Progressive sparsity regularization
@@
0:  Accumulated step numbers of each stage $\{T_{i}\}_{i=1}^{S}$, s.t. $0<T_{1}<T_{2}<...<T_{S}$.
@@
4:     update model by loss $\mathcal{L}_{lm}+\mathcal{L}_{reg}(\lambda)$
5:  end for
6:  // incremental stages
7:  for $i\leftarrow 2$ to $S$ do
8:     for $t\leftarrow T_{i-1}+1$ to $T_{i}$ do
9:        $\eta\leftarrow\frac{1}{2}[\sin(-\frac{\pi}{2}+\frac{t-T_{i-1}}{T_{i}-T_{i-1}}\pi)+1]$
10:        $\lambda\leftarrow\lambda_{i-1}+\eta(\lambda_{i}-\lambda_{i-1})$
11:        update model by loss $\mathcal{L}_{lm}+\mathcal{L}_{reg}(\lambda)$
12:     end for
13:  end for
@@
As demonstrated by recent works, there exist considerable amounts of non-zero low elements in the activation outputs, which have little influence on final results and thus can be pruned for higher sparsity [88]. Therefore, we convert the ReLU into FATReLU [38] by shifting the activation threshold, i.e.,
@@
where $t>0$ is a positive threshold. As long as $t$ is properly chosen (see Appendix J), FATReLU can increase sparsity with negligible losses [88].
3.3 Practical Inference Acceleration
To go further beyond former theoretical acceleration analyses based on FLOPS (Floating Point Operations Per Second) [50] and establish the practical value of ProSparse, in this section, we discuss how to realize inference acceleration with sparsely activated LLMs on real GPU hardware and how to evaluate the practical acceleration effects. We consider two categories of acceleration algorithms based on activation sparsity: the approximate algorithm and the accurate algorithm.